I was told to Repost this question
Evaluate the exponential equation for three positive values of x, three negative values of x, and at x=0. Show your work. Use the resulting ordered pairs to plot the graph; submit the graph via the Dropbox. State the equation of the line asymptotic to the graph (if any).
y=3^x-4+1
The exponent if I understand question is x-4

Answer by richwmiller(17219)

So now choose three positive value for x and three negative values for x and x=0
and solve for y
x=6 :y=3^(6-4)+1:y=3^(2)+1=9+1
x=5 :y=3^(5-4)+1=y=3^1+1=3+1
x=4 :y=3^(4-4)+1:y=3^0+1=1+1
x=-1:y=3(-1-4)+1:y=3^(-5)+1:y=1/3^5+1
x=-2:y=3(-2-4)+1:y=3^(-6)+1:y=1/3^6+1
x=-3:y=3(-3-4)+1:y=3^(-7)+1::y=1/3^7+1
x=0: y=3^(-4)+1:y=1/3^4+1
finish calculating and
plot those values